You are ordering a hamburger and can get up to 6 toppings, but each topping can only be used once. You tell the cashier to surprise you with the toppings you get. What is the probability that you get 2 toppings? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

Solution

Given that ,

p = 1 / 6 = 0.1667

1 - p = 1 - 0.1667 = 0.8333

n = 6

x = 2

Using binomial probability formula ,

P(X = x) = ((n! / x! (n - x)!) * px * (1 - p)n - x

P(X = 2) = ((6! / 2! (6 - 2)!) * 0.16672 * (0.8333)6 - 2

=   ((6! / 2! (4)!) * 0.16672 * (0.8333)4

= 0.2010

Probability = 0.2010